If you have 2.2 L of helium gas (He) originally at 15 °C and at 775 mmHg of pressure, what is the pressure of helium gas in mmHg if the final temperature is 75 °C and the volume is 1.5 L?
1
Expert's answer
2020-11-22T12:05:08-0500

V1 = 2.2 L

T1 = 15 +273.15 = 288.15 K

P1 = 775 mmHg

T2 = 75 + 273.15 = 348.15 K

V2 = 1.5 L

"\frac{P_1V_1}{T_1} = \frac{P_2V_2}{T_2}"

"P_2 = \frac{P_1V_1T_2}{T_1V_2}"

"P_2 = \frac{775 \times 2.2 \times 348.15}{288.15 \times 1.5} = 1373 \;mmHg"